1. Your current salary is $100,000 per year. Every year you save 10 percent of your salary and invest it in a retirement plan which provides you a guaranteed return of 6 percent per year. Assume the deposits will be made at the end of each year for the next 40 years.

a) How much will you have at the end of 40 years if your salary grows at 6 percent per year?

b) What annual fixed amount will you be able to withdraw every year for 20 years following retirement? (Assume you continue to earn 6 percent per year)
Answer #1

a) Rate of Return =6%
Number of years =40
growth =6%
PMT =Current Salary*Growth =100000*10% =10000
FV of amount =10000/(1+6%)+10000*(1+6%)/(1+6%)^2...........+10000*(1+6%)^39/(1+6%)^40
=10000/(1+6%)+10000/(1+6%)...........+10000/(1+6%)
=10000/(1+6%)*40=377358.49

b. Number of years =20
Rate =6%
Fixed Amount per year on retirement using annuity formula=PV/((1-(1+r)^-n)/r) =377358.49/((1-(1+6%)^-20)/6%) =32899.83